OpenStudy (anonymous):

what is the wavelength of a wave with a frequency of 4.00x10^8

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

You can figure this out from the following relationship: the wavelength multiplied by the frequency is equal to the speed of light: to calculate substitute 4e8 in and divide by 3e8 (speed of light in metres per second) and you will have your answer in metres. There's more explanation here: http://en.wikipedia.org/wiki/Planck%27s_constant

OpenStudy (mayankdevnani):

frequency = speed of light / wavelength so 4*10^8 = 3* 10^8 / wavelength so wavelength = ?

OpenStudy (mayankdevnani):

wavelength = 3*10^8 / 4*10^8 = ?

OpenStudy (mayankdevnani):

so your wavelength is 0.75 meters

OpenStudy (mayankdevnani):

@annej

OpenStudy (anonymous):

I stand corrected - mayankdevani is correct and I got my algebra wrong - I should have said "divide the speed of light by the wavelength".

OpenStudy (mayankdevnani):

oh!!!!